OpenStudy (anonymous):

Solve the following equation: (x^4 + 1)(4 + x) = 0

OpenStudy (anonymous):

u might use zero product property

OpenStudy (anonymous):

x^4+1=0 or 4+x=0

OpenStudy (anonymous):

4+x=0 gives x=-4 how about the other one?